OpenStudy (anonymous):

PLS HELP URGENNT MUST GET 100% on study guide!!!! If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x + y + 2z = 1 x + 2y – z = –8 3x – y – 3z = 18

OpenStudy (mathmate):

2(2)+(1) : 5x+5y = -15 => x+y=-3 3(2)-(3) : 7y = -42 => y = -6

OpenStudy (anonymous):

These are the answer choices. so it would be the 3rd answer? 7y – z = 25 5y – 6z = 42 –5y + 5z = 25 –7y = 42 –5x – 5z = –10 6x – z = 19 7x + 3z = –10 0x – 5z = 19

OpenStudy (mathmate):

Sorry, I eliminated z, so none of the equations match. Here I have eliminated x, with the following: 5y-5z = -25 7y = -42 So that's (B), I realize the signs have been inverted, but that depends on the order of subtraction. I like to subtract in the direction that gives a positive sign to the variables.